OpenStudy (anonymous):

\[\left(\begin{matrix}4c^2 \\4c^2-24c+36\end{matrix}\right)*\left(\begin{matrix}4c-12 \\ 2c\end{matrix}\right)\]

OpenStudy (anonymous):

Multlipy and simplify

hero (hero):

It's much simpler if you just use frac {num} {den} * frac {num} {den}

OpenStudy (anonymous):

Oh ok sorry

OpenStudy (anonymous):

Try this instead: ~[ \frac {4c^2} {4c^2-24c+36} \times \frac {4c-12} {2c} ~] And change every ~ to a \

OpenStudy (anonymous):

\[\frac {4c^2} {4c^2-24c+36} * \frac {4c-12} {2c }\]

hero (hero):

It automatically formatted it for him.

OpenStudy (anonymous):

There ya go. Now the first thing you'll want to do is factor the denominator of that first fraction.

OpenStudy (anonymous):

Sorry everyone, thanks

myininaya (myininaya):

2*1 times a 2*1 so such multiplication can be performed

myininaya (myininaya):

no*

myininaya (myininaya):

wait is he writing fractions or matrices?

myininaya (myininaya):

it looks like matrices

OpenStudy (anonymous):

They are fractions

OpenStudy (zarkon):

disappointed..thought this was going to be some cool combinatoric problem ;)

OpenStudy (anonymous):

lol, yes, just multiplication of fractions guys.

OpenStudy (anonymous):

It's for algebra 2

myininaya (myininaya):

oops earlier i thought he was adding matrices not fractions so i guess he got the wrong answer for that one question earlier

OpenStudy (anonymous):

Sorry again everyone I'm a noob and horrible with math! :(

hero (hero):

Don't put yourself down so easily there bud

OpenStudy (anonymous):

A lot of people like to say they're bad at math, even if they aren't. I don't really understand the group-think behind that.

OpenStudy (anonymous):

lol, thanks to clear the air I am a female, LMAO

OpenStudy (anonymous):

So are you still working on this problem? Did you factor the denominator?

OpenStudy (anonymous):

yes I got 4c^2+48c

OpenStudy (anonymous):

females are good at math.

OpenStudy (anonymous):

No, that's not quite right.

OpenStudy (anonymous):

LOL, oh not this one!!!!

OpenStudy (anonymous):

did I forget the 36

OpenStudy (anonymous):

4c^2+48c+36 Is that right?

OpenStudy (anonymous):

Bah, that's wrong actually. It should be 4(c + a)(c+b) So we need to factor a 4 from everything first.

OpenStudy (anonymous):

Can you factor a 4 from each term?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

I would divide them by 4 right?

myininaya (myininaya):

i'm a female too so if i can do this you can

OpenStudy (anonymous):

Bringing the 4 out in front leaves 1 less factor of 4 in each, so yes, you divide by 4 to each of them to see what's left.

myininaya (myininaya):

don't be sexist against your own kind lol

OpenStudy (anonymous):

c^2-6c+9?

OpenStudy (anonymous):

With a 4 out in front, and that part in parens, yes.

OpenStudy (anonymous):

\[4c^2 - 24c + 36 = 4(c^2 - 6c + 9)\]

OpenStudy (anonymous):

OH ok

OpenStudy (anonymous):

Would I do the same for the top?

OpenStudy (anonymous):

Myin, she can do this. Let her. =)

myininaya (myininaya):

polpak i didn't do the whole thing lol i was gonna let her finish the rest of it

OpenStudy (anonymous):

Yeah, but this part is the hardest part. Don't steal the fun ;)

OpenStudy (anonymous):

lol, thank you for the confidence everyone :)

OpenStudy (anonymous):

ok so this is what I got for the top 16c^3-12?

OpenStudy (anonymous):

wait huh?

OpenStudy (anonymous):

The top of what?

OpenStudy (anonymous):

The first fraction or the second?

OpenStudy (anonymous):

Don't do the multiplication of the two fractions until after you simplify them, otherwise it'll mean more work later.

OpenStudy (anonymous):

The fraction, the 4c^2 and 4c-12

OpenStudy (anonymous):

ok now I am confused, I'm sorry

OpenStudy (anonymous):

It's ok. You see that there are two fractions here being multiplied right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok to clear this up for me, the part of the problem we just did was just the denominator for the first fraction right?

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

We factored it. Now since there is a factor of 4 on top and bottom you can cancel it right?

OpenStudy (anonymous):

\[\frac {4c^2}{4c^2-24c+36} * \frac{4c-12}{2c }\]\[=\frac {4c^2} {4(c^2-6c+9)} * \frac {4c-12} {2c }\]

OpenStudy (anonymous):

This is what we've done so far.

OpenStudy (anonymous):

Right, so would that cancel out the 4c^2 on top?

OpenStudy (anonymous):

You cannot cancel the c^2 because we don't have a factor of c^2 on bottom. But we do have a factor of 4 on each, so the 4's (only) cancel.

OpenStudy (anonymous):

Factors are things being multiplied to make a product. The c^2 you have on the bottom is one of the terms (somthing being added) to make a factor, but it is not a factor by itself.

OpenStudy (anonymous):

ok, so then would it be \[\left(\begin{matrix}c^2 \\ c^2-62+9\end{matrix}\right)\]

OpenStudy (anonymous):

Sorry I keep putting it in wrong

OpenStudy (anonymous):

That's ok, you can use combinations for now, I'll know what you mean ;)

OpenStudy (anonymous):

That 2 on the bottom should be a c right?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

lol, yes sorry

OpenStudy (anonymous):

Ok, good. Now then, lets look at the other fraction: \[=\frac {4c^2} {4(c^2-6c+9)} * \frac {4c-12} {2c }\] We have a 2 on bottom, can we factor out a 2 from the top?

OpenStudy (anonymous):

Whoops, forgot to cancel the 4's you got rid of: \[=\frac {c^2} {c^2-6c+9} * \frac {4c-12} {2c }\]

OpenStudy (anonymous):

So can you factor a 2 from the top of the right hand fraction?

OpenStudy (anonymous):

So would the other fraction be \[2c-12\]

OpenStudy (anonymous):

you have to factor a 2 from both terms in the numerator, you cannot cancel a 2 from just part of the factor.

OpenStudy (anonymous):

4c-12 is a factor by itself, you have to extract from it a factor of 2. To do that (just like the denominator of the other fraction) we have to pull a 2 from each term, 4c and -12.

OpenStudy (anonymous):

Then we can cancel the factor of 2 from the top and the bottom.

OpenStudy (anonymous):

ok so 2c-6

OpenStudy (anonymous):

That's the numerator after you cancel, what is left in the denominator?

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

Also correct.

OpenStudy (anonymous):

really?!

OpenStudy (anonymous):

\[=\frac {c^2} {c^2-6c+9} * \frac {4c-12} {2c }\]\[=\frac {c^2} {c^2-6c+9} * \frac {c-6} {c }\] Yes really. I told you that you weren't bad at math ;p

OpenStudy (anonymous):

Thank you so so much!!!! I'm so excited!

OpenStudy (anonymous):

So now lets combine the two, but just as factors, don't multiply them all out yet.

OpenStudy (anonymous):

When you multiply two fractions, what's the proceedure?

OpenStudy (anonymous):

Just describe it, don't try to give the result yet.

OpenStudy (anonymous):

um, numerator by denominator (multiplying)

OpenStudy (anonymous):

Right!, So lets just rewrite them like that without doing the whole distributing thing.. Just put them together with some parens.. What do you have on the top and what on the bottom?

OpenStudy (anonymous):

On the top c^2(c)? and the bottom c^2-6c+9(c-6)?

OpenStudy (anonymous):

That doesn't look right to me

OpenStudy (anonymous):

No, you flipped the right hand fraction for some reason.

OpenStudy (anonymous):

\[=\frac {c^2} {c^2-6c+9} * \frac {2c-6} {c }\]

OpenStudy (anonymous):

and you lost the 2 in front of the c on the c-6 also

OpenStudy (anonymous):

Ok, so should it be, top c^2(c) bottom 2c-6(c^2-6c+9)?

OpenStudy (anonymous):

No. You still flipped them. Don't do that. Just multiply straight across.

OpenStudy (anonymous):

Also put parens around the 2c-6, otherwise it looks like you just multipling the -6.

OpenStudy (anonymous):

\[\frac{2}{5}\times \frac{3}{7} = \frac{2(3)}{5(7)}\]

OpenStudy (anonymous):

c^2(2c-6) c^2-6c+9(c)

OpenStudy (anonymous):

That's right on top. On bottom you put parens around the wrong thing again. If the denominator has multiple terms, put parens around it when you multiply.